OpenStudy (anonymous):

Rationalize each denominator and simplify 3/√7 +5

OpenStudy (saifoo.khan):

\[\frac{5-\sqrt7}{6}\]

OpenStudy (anonymous):

may i also ask your process of getting to that answer?

OpenStudy (anonymous):

Assuming that the square root affects all of the denominator, the easiest way to simplify is by using the conjugate method. Multiply by \[\sqrt{7-5}/\sqrt{7-5}\] and simplify.

OpenStudy (saifoo.khan):

\[\frac{3}{\sqrt7 + 5} \times \frac{\sqrt7 - 5}{\sqrt7 - 5}\]